The weight W that a horizontal beam can support varies inversely as the length L of the beam. Suppose an 8 in beam can support 1200 kg. How many kilograms can a 14-m beam support?
Answer by Alan3354(69443)

The weight W that a horizontal beam can support varies inversely as the length L of the beam. Suppose an 8 in beam can support 1200 kg. How many kilograms can a 14-m beam support?
-------------
W = k/L
W*L = k
---------
8 in = 0.2032 meters
----
0.2032*1200 = 14*W
W = 0.2032*1200/14
W =~ 17.417 Kg